OpenStudy (anonymous):

Use exponents properties to simplify (7^1/4)(7^1/2)

OpenStudy (anonymous):

Can you explain the steps

OpenStudy (anonymous):

your job is to add \(\frac{1}{2}+\frac{1}{4}\)

OpenStudy (anonymous):

because \(b^n\times b^m=b^{m+n}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

I got 1/2+1/4=2/2=1

OpenStudy (anonymous):

=49